Wednesday, March 25, 2020, 2:00 pm EDT

From Hall A Wiki
Jump to: navigation, search

Back to Main Compton Page >> Compton Meetings

*********BLUEJEANS CALL-IN INSTRUCTIONS*********
----------------------------------------------------
1. Dial Toll-Free Number: 1-888-240-2560 (U.S. & Canada)
2. International participants dial:
Toll Number: 1-408-740-7256
3. Enter meeting ID:140287970

Desktop sharing:
https://bluejeans.com/140287970 

Agenda

Meeting room = None. Remote only.

General

  • Compton status after shutdown
  • Compton DAQ/controls racks will probably move!

Laser

  • Laser polarization systematic tests [1]

Photon Detector

  • Beam steering test results [2]
  • Linearity measurements

Analysis

  • Online analysis results
  • Polarization vs. time
  • Offline analysis plans

Electron detector

  • Analysis updates


Minutes

Dave G, Adam Z, Amali, Cip G, Don J, JCC, Mark D, Kent P, Alex C, Brian Q

  1. General
    1. Due to SBS, Compton racks will move to place unknown.
      1. Need to document everything so that it can be reestablished
  2. Laser
    1. Online results show consistent results for the alternate polarization state.
  3. Photon Detector
    1. Spectrum
      1. Managed to get a measurable offset in the spectrum.
      2. When we project the BPMs to the collimator, then y=0mm is not the centre. It looks like y=+1mm is the centre.
      3. We need a summary of all the pointing spectrum runs.
    2. Linearity
      1. Variable LED appears to have broken. Couldn't be repaired in the time available.
      2. More linearity runs needed to give equivalent load to the CREX running conditions.
  4. Analysis
    1. Begin offline analysis with a respin, particularly of early data with improved cycle detection.